A food company distributes its tomato soup in two cans of different sizes. For the larger can, both the diameter and the height have been increased by 10%. By what percentage does the volume of the can increase from the smaller can to the larger can? Round your answer to the nearest percent.
Answer #1

Answer :

The volume of a can with radius r and height h is

V = πr2h

when the original can with r = 1 and h = 1 , we have

V = π(1)2(1) = π

When both diameter and height increased by 10% :

That is when r = 1.1 and h = 1.1 , we have V = π(1.1)2(1.1) = 1.331π

The increase in volume is 0.331π

Now Consider (0.331π/π)* 100% = 0.331*100 = 33.1 %

The percentage increase is 33.1%